A particle's position on the x axis is given by x=5t^3+20t^2-t+10





Fine the particle velocity function v(1)and acceleration a (t)

1
Expert's answer
2022-03-13T18:46:29-0400

Velocity is the first derivative of the position:


"v(t) = 15t^2 + 40t + 1\\
v(1) = 15\cdot 1^2 + 40\cdot 1 + 1 = 56"

Acceleration is the second derivative of the position:


"a(t) = 30t + 40"